Q8. Calculate the divergence and Curl, of the following functions

a. V= x2 i+3xy^2 j -2xyz k

b. V=xyi+2yz j+3zx k

c. V= y2i +(2xy+z2)j+ 2yz k , where i , j , k are unit vectors along x, y and z axis


1
Expert's answer
2021-09-18T15:00:23-0400
"{\rm div}{\bf V}=\frac{\partial V_x}{\partial x}+\frac{\partial V_y }{\partial y}+\frac{\partial V_z}{\partial z}"

"\rm curl\:{\bf V}=\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
V_x&V_y&V_z
\end{vmatrix}"


(a)

"{\rm div}{\bf V}=\frac{\partial (x^2)}{\partial x}+\frac{\partial (3xy^2) }{\partial y}+\frac{\partial (-2xyz)}{\partial z}\\=2x+6xy-2xy=2x+4xy"

"{\rm curl\:{\bf V}}=\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
x^2&3xy^2&-2xyz
\end{vmatrix}""=-2xy\hat i+2yz\hat j+3x\hat k"

(b)

"{\rm div}{\bf V}=\frac{\partial (xy)}{\partial x}+\frac{\partial (2yz) }{\partial y}+\frac{\partial (3zx)}{\partial z}\\=y+2z+3x"

"{\rm curl\:{\bf V}}=\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
xy&2yz&3zx
\end{vmatrix}"

"=-2y\hat i-3z\hat j-x\hat k"

(c)

"{\rm div}{\bf V}=\frac{\partial (y^2)}{\partial x}+\frac{\partial (2xy+z^2) }{\partial y}+\frac{\partial (2yz)}{\partial z}\\=2x+2y"

"{\rm curl\:{\bf V}}=\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
y^2&2xy+z^2&2yz
\end{vmatrix}"

"=0\hat i+0\hat j+0\hat k=\bf 0"